A ballon starts rising from a ground with constant accelation of 1.28.After 8sec a stone is released . find the time taken by the stone to reach the ground

Grade:9

1 Answers

Khushi goyal
6 years ago
Disp of the lift in 8 sec = 1/2×1.28×6 =40.96 (initial vel is 0 as it starts rising without any mentioned force )Velocity of lift after 8 sec =0+1.28×8=10.24After 8 sec when the stone is released it will have the same relative velocity.coz it is not given whether it is released from rest ..or something like that.So. 40.96=10.24t +1/2gt^2 Solve this quadratic and get t.If you got it then plz click approved. Good luck !